OpenStudy (jkbo):

PLEASE HELP What is the equation of the circle with center (2, –5) that passes through the point (–2, 10)? (1 point) (x – 2)2 + (y – (–5))2 = 25 (x – (–2))2 + (y – 10)2 = 241 (x – 2)2 + (y – (–5))2 = 241 (x – (–2))2 + (y – 10)2 = 25

OpenStudy (campbell_st):

find the distance from the centre (2, -5) to the point (-2, 10) which is on the circumference this will be the radius. The general form of the circle is \[(x -h)^2 + (y -k)^2 = r^2\] (h, k) is the centre and r is the radius.

OpenStudy (campbell_st):

so the partial equation is \[(x - 2)^2 + (y - (-5))^2 = r^2\]

OpenStudy (jkbo):

i need help solving it please

OpenStudy (jkbo):

@campbell_st

OpenStudy (campbell_st):

do you know the distance formula..?

OpenStudy (jkbo):

no i just dont get how to solve the equation

OpenStudy (jkbo):

ok i solved it and i got B. @campbell_st

OpenStudy (jkbo):

is that right?